OpenStudy (anonymous):

Can someone help me with Matrix Multiplication? I will post question in the comments. I go to Connexus if anyone can help that way.

OpenStudy (anonymous):

Find -5A+4B 6 1 A= -4 -6 7 -7 -5 -1 B= -3 -8 6 8

OpenStudy (anonymous):

@nincompoop Can you help?

OpenStudy (chaise):

Multiply every entry by -5 in matrix A, and then multiply every entry in matrix B by 4, then the new matrix you make, add them together.

OpenStudy (anonymous):

Okay thank you.

OpenStudy (chaise):

When you add matrices together, you add every entry with every other respective entry. For example: A=[1, 2] and B=[3,4] A+B=[4, 6]

OpenStudy (anonymous):

I knew how to add I just got confused with multiplication.

OpenStudy (anonymous):

Dont mix up dot product, cross product, and matrices times a scalar